OpenStudy (anonymous):

Solve the optimization problem. Minimize F = x2 + y2 with x + 2y = 5. F= ?

OpenStudy (anonymous):

\[F=(5-2y)^2+y^2\] \[F(y)=25-20y+5y^2\]

OpenStudy (anonymous):

\[F^'(y)=10y-20=0\] \[y=2\] \[x=5-2(2)=1\] \[F(1,2)=1^2+2^2=5\]

OpenStudy (anonymous):

Thanks for the help!!